# Bootstrap Off-policy with World Model

### v2

@@ 1 Introduction
Reinforcement learning (RL) has achieved impressive performance in a wide range of domains, from industrial automation to autonomous driving and embodied intelligence vinyals2019grandmaster; schrittwieser2020mastering; zhan2024transformation. Among the various techniques developed to enhance RL, online planning stands out for its predictive optimization ability to improve control performance using learned dynamics moerland2023model; campbell2019model; zhan2023enhance. By performing look-ahead rollouts, it enables agents to anticipate future consequences and iteratively refine actions tassa2012synthesis; schrittwieser2021online. Compared to model-free approaches, which rely solely on trial-and-error learning, model-based planning offers an effective tool to generate high-quality actions for environment interaction deisenroth2011pilco; curi2020efficient.
@@
A growing body of research has explored how to more effectively integrate planning into RL hamrickrole; moerland2023model. Early methods such as PETS chua2018deep and PlaNet hafner2019learning have proven that planning with learned dynamics and reward signals can achieve impressive control performance. To further improve performance in high-dimensional tasks, recent approaches have combined online planning with policy learning, where the policy can provide a good initial solution to speed up planning argenson2020model. For example, LOOP sikchi2022learning builds on SAC haarnoja2018soft, an off-policy model-free algorithm, and integrates a planner guided by a learned dynamics model to enable higher-quality environment interaction. Recently, TD-MPC and TD-MPC2 hansen2022temporal; hansen2023td jointly learn the dynamics, reward and value functions through temporal differential (TD) learning, achieving strong performance through both algorithmic innovations and implementation advances. These methods successfully deliver substantial gains over model-free baselines such as SAC, particularly on high-dimensional benchmarks.
@@
However, this class of planning-driven model-based RL algorithms inevitably suffers from a fundamental issue known as actor divergence: the data used for learning is collected by the planner, which acts as a different actor from the policy network sikchi2022learning. Under the paradigm framework of off-policy RL, this issue leads to two problems: (1) Distribution shift in value learning: The value function is trained on the data collected by the planner rather than the policy itself. As a result, it learns accurately within the planner’s state-action distribution but tends to overestimate values in out-of-distribution regions that are rarely visited kumar2020conservative. (2) Unreliable policy improvement: The policy network is updated using value estimates from the value network, which are influenced by the distributional shift. These biased estimates may mislead the policy, impairing its ability to distinguish between good and bad actions, ultimately leading to severe performance degradation prudencio2023survey. To conclude, the misalignment between the training data and the policy’s actual behavior can severely hinder the learning process, particularly in complex, high-dimensional environments where accurate value estimation is already challenging nauman2024overestimation. The resulting value bias not only misguides policy updates but also risks destabilizing the entire training process markowitz2024avoiding. More critically, since online planning algorithms typically rely on sample-based optimization, the resulting action distributions are non-parametric and difficult to access. This means that they cannot be explicitly represented and the likelihood is intractable williams2017model.
@@
In this paper, we introduce three key contributions: (1) To facilitate alignment with the online planner’s sample-based non-parametric distribution, we adopt a likelihood-free alignment loss that measures the divergence between the policy and the planner without requiring explicit likelihoods of actions from the online planner. (2) We introduce a soft value-weighted mechanism that prioritizes high-return behaviors, driven by the planner’s value-guided action selection principle. Additionally, to maintain training efficiency, we align the policy with the stored planner actions in the replay buffer. Since these historical actions may vary in quality, our value-weighting mechanism ensures the policy prioritizes the high-valued promising experiences, accelerating learning while handling the variability in the planner’s past actions. (3) BOOM combines ease of implementation with state-of-the-art (SOTA) performance on high-dimensional continuous control benchmarks, including the DeepMind Control Suite tassa2018deepmind and Humanoid-Bench sferrazza2024humanoidbench.
@@ 2.1 Reinforcement Learning
Reinforcement learning (RL) offers a powerful framework for sequential decision-making li2023rlbook, formalized as a Markov Decision Process (MDP) $(\mathcal{S},\mathcal{A},P,r,\gamma)$, where $\mathcal{S}$ and $\mathcal{A}$ denote the state and action spaces, $P(s^{\prime}|s,a)$ is the transition dynamics, $r(s,a)$ is the reward function, and $\gamma\in[0,1)$ is the discount factor. At each timestep, an agent observes state $s_{t}$, selects action $a_{t}$, receives reward $r_{t}$, and transitions to the next state $s_{t+1}$ according to $P$. A fundamental concept in RL is the action-value function: $Q^{\pi}(s,a)=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\,\big|\,s_{0}=s,a_{0}=a\right],$ which captures the expected return of taking action $a$ in state $s$, followed by policy $\pi$. Off-policy RL aims to discover the optimal policy $\pi^{*}$ by maximizing the Q-value function using transitions generated from a different behavior policy $\beta$. By decoupling data collection from policy improvement, off-policy methods can achieve high sample efficiency—primarily due to the effective reuse of past transitions stored in a replay buffer. However, when the behavior policy $\beta$ deviates too far from $\pi$, the resulting distributional shift can severely undermine value estimation, a delicate yet critical challenge that continues to motivate much of the recent progress in off-policy learning DSACT.
@@ 2.2 Online Planner
Online planning optimizes action sequences at each step by simulating future trajectories under a predictive model, enabling informed action selection zhan2024explicit. This approach typically yields high-quality decisions by leveraging foresight over potential future outcomes. As a result, it is widely adopted in model-based RL for environment interaction, generating higher-quality interaction samples that facilitate more efficient learning and policy improvement.
Among various planning methods, Model Predictive Path Integral (MPPI) is a widely adopted sampling-based planner for continuous control tasks due to its high efficiency williams2017model. Formally, at each planning step, with the help of a dynamics model, $N$ action trajectories are sampled from a factorized Gaussian: $a_{t}^{i}\sim\mathcal{N}(\mu_{t},\sigma_{t}^{2}),\quad\text{for }t=0,\dots,H{-}1\text{ and }i=1,\dots,N,$ where $H$ is the predictive horizon. Each trajectory is evaluated using a reward and value model to estimate the future return: $G^{i}=\sum_{t=0}^{H-1}r_{t}^{i}+\hat{v}_{H}^{i},$ where $r_{t}^{i}$ denotes predicted rewards and $\hat{v}_{H}^{i}$ is the terminal value estimate. To update the trajectory distribution, MPPI reweights the $N$ samples using a softmax over the returns: $w^{i}={\exp(G^{i}-\max_{j}G^{j})}/{\sum_{k}\exp(G^{k}-\max_{j}G^{j})}.$ Next, the parameters of the trajectory distribution are updated using weighted statistics: $\mu_{t}\leftarrow\sum_{i}w^{i}a_{t}^{i}.$ This process is repeated over several iterations, gradually converging toward an optimal solution. At test time, the planner executes the first action $\mu_{0}$, while during training, Gaussian noise is added for exploration.
@@ 3.2 Bootstrap Off-policy with World Model
$\mathcal{L}_{\text{align}}=\mathbb{E}_{(s,a,r,s^{\prime})_{0:H}}\sum_{i=1}^{N}w_{i}\left[-\log\pi(a_{i}\mid s_{i})\right].$ (4)
This mechanism ensures that the policy assigns higher probability to high-value actions, effectively guiding it toward promising regions. Besides, since these actions stored in the replay buffer may vary in quality, this value-weighting ensures that the policy prioritizes the most beneficial experiences, thus accelerating learning while accommodating the variability in the planner’s past actions. The Q function here can also be replaced by other critics, such as state value or advantage functions, but we think Q is the most convenient one to access.
@@
$\mathcal{L}_{\text{policy}}=-\mathbb{E}_{(s,a,r,s^{\prime})_{0:H}}\sum_{t=0}^{H-1}Q(s,\pi(s))+\lambda_{\text{align}}\cdot\mathcal{L}_{\text{align}},$ (5)
@@
The reasons are two fold. (1) Improve value learning. By aligning the policy with the planner, we reduce the discrepancy between the collected data and the actual policy behavior. This improves the distributional matching during training, enabling the value estimator to learn from more consistent and policy-relevant trajectories. As a result, the predicted values used in planning become more accurate and reliable. (2) Improve representation, reward, and dynamics learning. We adopt a TD-style learning objective that jointly optimizes the value function along with the other components of the world model as shown in (1). As the value predictions become more accurate due to better distributional match, the gradients flowing into the encoder, dynamics model, and reward predictor become more informative, leading to improved overall model quality.
@@
12:  Replay rollout batch $\{(s_{t},a_{\beta},r_{t},s_{t+1})_{t=0}^{H}\}\sim\mathcal{D}$
@@ 4.1 Experimental Setup
We choose four representative online RL algorithms as our baselines: (1) SAC haarnoja2018soft: the state-of-the-art model-free off-policy RL algorithm under the maximum entropy framework; (2) DreamerV3 hafner2025mastering: the state-of-the-art imagination-driven model-based RL algorithm that learns from imaginary rollouts generated by a world model. To enable a comprehensive comparison, we evaluate DreamerV3 under two different interaction budgets: 4M and 10M environment steps; (3) TD-MPC2 hansen2023td: a planning-driven model-based RL algorithm that performs online planning via an MPPI planner; (4) BMPC wang2025bootstrapped: an improved variant of TD-MPC2, where the policy is trained solely to imitate the actions generated by the MPPI with an extra relabeling mechanism to update the historical actions in the replay buffer with the latest planner.
@@
We evaluate our method on a challenging benchmark of 14 high-dimensional locomotion tasks drawn from the DeepMind Control Suite (DMC) tassa2018deepmind and the recently proposed Humanoid Bench (H-Bench) sferrazza2024humanoidbench. The chosen 7 DMC tasks feature two most complex agents—humanoid (67/24 state/action dims) and dog (223/38)—that demand sophisticated balance and coordination. The other 7 H-Bench tasks raise the difficulty further with long-horizon, goal-directed tasks on the Unitree H1hand robot (151/61), such as walking over slides, traveling over a pole forest without collision, and continuously crossing hurdles. Detailed descriptions are listed in Appendix B.1.
Implementation details. We build upon the official TD-MPC2 codebase with our newly proposed Q-weighted likelihood-free bootstrap alignment. The detailed hyperparameters and reproducibility statement of other baselines are documented in Appendix B.2.
@@ 4.2 Experimental Results
All the training curves are shown in Figure 2 and the detailed numerical results are listed in Table 1. Additional training curves, including those under the 10M environment step setting, are provided in Appendix C.3. Our method, BOOM, consistently delivers the best Total Average Return (TAR) across all 14 high-dimensional locomotion tasks. These tasks pose significant challenges due to their large state and action spaces, yet BOOM demonstrates remarkable stability and effectiveness.
Results on the DMC Suite
. Our BOOM achieves an average TAR of 877.7, outperforming the previous best BMPC (835.8) by a substantial +5.0%, and exceeding TD-MPC2 by an even larger margin of +17.7%. Notably, BOOM achieves new best results on all humanoid and dog tasks. In Humanoid-run, BOOM outperforms the second-best method by +9.7%, and in Dog-run, it leads by a staggering +21.8%. While SAC and DreamerV3 often fail on these high-dimensional control tasks and achieve near-zero performance, planning-based methods like TD-MPC2 and BMPC perform better but still suffer from instability and limited final returns. In contrast, BOOM learns faster and achieves much stronger final performance across all tasks.
Results on the Humanoid Bench
. Our BOOM again dominates with an average TAR of 820.6. This marks a dramatic +47.7% improvement over DreamerV3 (10M), which is (555.6), and an even more impressive +60.5% gain over BMPC. BOOM sets new records on every single task in the Humanoid Bench. For example, in H1hand-slide, BOOM improves over the second-best method by +110.5%, in H1hand-pole, by +25.8%, and in H1hand-hurdle, by +121.0%, .
@@
Task SAC DreamerV3 ( 4M & 10M steps) TD-MPC2 BMPC BOOM (ours)
@@ 4.3 Ablation Study
We vary the alignment coefficient by testing $0.1\times$ and $10\times$ the default setting ($\lambda_{\text{align}}=\dim(\mathcal{A})/1000$ for DMC and $\dim(\mathcal{A})/100$ for Humanoid-Bench). Results in Figure 3(c) show stable performance across this range, suggesting that BOOM is robust to this hyperparameter and does not require very sensitive tuning.
@@ 5 Related Work
Planning-driven methods use planner rather than policy itself to generate high-quality actions for environment interaction curi2020efficient. Early work has shown that solely learning the value and then combining it with the planner can achieve good control performance  nagabandi2018neural; kocijan2004gaussian; deisenroth2013gaussian; lowreyplan. To further improve performance in high-dimensional tasks, recent approaches have combined online planning with policy learning, where the policy can provide a good initial solution to speed up planningjanner2019trust; wang2019exploring; nguyen2021temporal; buckman2018sample; lin2025td. LOOP sikchi2022learning takes SAC haarnoja2018soft as the backbone, and employs a planner under the policy behavior constraint for collecting samples. TD-MPC family hansen2022temporal; hansen2023td jointly learns model and value function through TD-learning, achieving strong performance through both algorithmic innovations and implementation advances. It successfully delivers substantial gains over model-free baselines, particularly on complex benchmarks. However, these methods inevitably encounter actor divergence—a mismatch between the planner and the policy. Our approach, BOOM, addresses this challenge by tightly coupling planning and off-policy learning through a bootstrap loop that aligns the policy with the planner’s non-parametric action distribution via a Q-weighted likelihood-free alignment loss, preserving distributional consistency. A recent method BMPC wang2025bootstrapped simplifies the pipeline by discarding explicit policy optimization and directly imitating the planner to avoid actor divergence. However, BMPC ignores the Q-function during training, resulting in lower policy learning efficiency and sensitivity to the variability of historical planner actions in the buffer. This leads to unstable learning, as reflected by its oscillatory training curves in our experiments. We acknowledge a concurrent and close work, TDM(PC)2 lin2025td, which similarly found that aligning the policy with the planner is beneficial. The major distinction lies in the design of the alignment objective: their approach follows a TD3+BC style using reverse KL, whereas ours is closer to AWAC using critic-guided weights and forward KL.
@@
Imagination-driven methods leverage a learned model to generate synthetic rollouts for policy and value updates sutton1991dyna. Modern approaches such as SimPLe kaiser2019model, IRIS micheli2022transformers, IDM mu2021model, and the Dreamer family hafner2019dream; hafner2020mastering; hafner2023mastering train latent dynamics models to support actor-critic learning entirely in imagination. These methods offer fast test-time execution and relatively high sample efficiency, but their performance is often limited by compounding model errors over long imagined rollouts. Among them, DreamerV3 hafner2025mastering stands out as a leading representative, demonstrating strong performance across a range of tasks. Unlike planning-based MBRL methods, DreamerV3 interacts by directly sampling actions from its learned policy, typically requiring more environment steps to converge. For comprehensive comparison, we evaluate DreamerV3 under 4M and 10M step budgets. While it improves with more interaction budget and outperforms TD-MPC2 and BMPC on certain tasks, our BOOM consistently outperforms all baselines across all tasks.
@@
One might notice the similarity between BOOM and offline RL; however, the fundamental paradigms and practical implications of these two settings differ significantly. (1) Learning paradigm. Offline RL centers on a fixed dataset, relying heavily on behavior cloning (BC) to restrict the policy within the dataset’s support fujimoto2019off. To cautiously improve policy performance, offline RL relaxes BC for poor actions, but always aims to keep the policy close to known data to avoid extrapolation error prudencio2023survey; ran2023policy. BOOM centers on the policy itself as the optimization target. It seeks to maximize Q-values and align with planner-generated actions, both aimed at improving policy performance. Ultimately, the policy and planner co-adapt and converge to the optimal solution through online interaction. (2) Learning objective. In offline RL, maximizing Q-values and BC often conflict fujimoto2021minimalist: Q-values outside the dataset distribution cannot be accurately estimated, so maximizing Q risks pushing the policy toward unsupported actions; BC pulls the policy back toward known actions. This tension forces a conservative balance wang2022diffusion. In BOOM, the two objectives are largely complementary: the planner generally produces higher-quality actions than the policy zhan2025bicriteria. Aligning policy with the planner improves performance and strengthens policy-planner consistency, which in turn leads to more accurate Q estimates. More accurate Q-values then enable better policy improvement and allow the planner to generate higher-quality actions. This positive feedback loop bootstraps policy and planner consistently toward faster convergence to the optimal solution.

### v3

@@ 1 Introduction
Reinforcement learning (RL) has achieved impressive performance in a wide range of domains, from industrial automation to autonomous driving and embodied intelligence Vinyals et al. (2019); Schrittwieser et al. (2020); Zhan et al. (2024b). Among the various techniques developed to enhance RL, online planning stands out for its predictive optimization ability to improve control performance using learned dynamics Moerland et al. (2023); Campbell et al. (2019); Zhan et al. (2023). By performing look-ahead rollouts, it enables agents to anticipate future consequences and iteratively refine actions Tassa et al. (2012); Schrittwieser et al. (2021). Compared to model-free approaches, which rely solely on trial-and-error learning, model-based planning offers an effective tool to generate high-quality actions for environment interaction Deisenroth and Rasmussen (2011); Curi et al. (2020).
@@
A growing body of research has explored how to more effectively integrate planning into RL Hamrick et al. (2020); Moerland et al. (2023). Early methods such as PETS Chua et al. (2018) and PlaNet Hafner et al. (2019b) have proven that planning with learned dynamics and reward signals can achieve impressive control performance. To further improve performance in high-dimensional tasks, recent approaches have combined online planning with policy learning, where the policy can provide a good initial solution to speed up planning Argenson and Dulac-Arnold (2020). For example, LOOP Sikchi et al. (2022) builds on SAC Haarnoja et al. (2018), an off-policy model-free algorithm, and integrates a planner guided by a learned dynamics model to enable higher-quality environment interaction. Recently, TD-MPC and TD-MPC2 Hansen et al. (2022, 2023) jointly learn the dynamics, reward and value functions through temporal differential (TD) learning, achieving strong performance through both algorithmic innovations and implementation advances. These methods successfully deliver substantial gains over model-free baselines such as SAC, particularly on high-dimensional benchmarks.
@@
However, this class of planning-driven model-based RL algorithms inevitably suffers from a fundamental issue known as actor divergence: the data used for learning is collected by the planner, which acts as a different actor from the policy network Sikchi et al. (2022). Under the paradigm framework of off-policy RL, this issue leads to two problems: (1) Distribution shift in value learning: The value function is trained on the data collected by the planner rather than the policy itself. As a result, it learns accurately within the planner’s state-action distribution but tends to overestimate values in out-of-distribution regions that are rarely visited Kumar et al. (2020). (2) Unreliable policy improvement: The policy network is updated using value estimates from the value network, which are influenced by the distributional shift. These biased estimates may mislead the policy, impairing its ability to distinguish between good and bad actions, ultimately leading to severe performance degradation Prudencio et al. (2023). To conclude, the misalignment between the training data and the policy’s actual behavior can severely hinder the learning process, particularly in complex, high-dimensional environments where accurate value estimation is already challenging Nauman et al. (2024). The resulting value bias not only misguides policy updates but also risks destabilizing the entire training process Markowitz et al. (2024). More critically, since online planning algorithms typically rely on sample-based optimization, the resulting action distributions are non-parametric and difficult to access. This means that they cannot be explicitly represented and the likelihood is intractable Williams et al. (2017).
@@
In this paper, we introduce three key contributions: (1) To facilitate alignment with the online planner’s sample-based non-parametric distribution, we adopt a likelihood-free alignment loss that measures the divergence between the policy and the planner without requiring explicit likelihoods of actions from the online planner. (2) We introduce a soft value-weighted mechanism that prioritizes high-return behaviors, driven by the planner’s value-guided action selection principle. Additionally, to maintain training efficiency, we align the policy with the stored planner actions in the replay buffer. Since these historical actions may vary in quality, our value-weighting mechanism ensures the policy prioritizes the high-valued promising experiences, accelerating learning while handling the variability in the planner’s past actions. (3) BOOM combines ease of implementation with state-of-the-art (SOTA) performance on high-dimensional continuous control benchmarks, including the DeepMind Control Suite Tassa et al. (2018) and Humanoid-Bench Sferrazza et al. (2024).
@@ 2.1 Reinforcement Learning
Reinforcement learning (RL) offers a powerful framework for sequential decision-making Li (2023), formalized as a Markov Decision Process (MDP) $(\mathcal{S},\mathcal{A},P,r,\gamma)$, where $\mathcal{S}$ and $\mathcal{A}$ denote the state and action spaces, $P(s^{\prime}|s,a)$ is the transition dynamics, $r(s,a)$ is the reward function, and $\gamma\in[0,1)$ is the discount factor. At each timestep, an agent observes state $s_{t}$, selects action $a_{t}$, receives reward $r_{t}$, and transitions to the next state $s_{t+1}$ according to $P$. A fundamental concept in RL is the action-value function: $Q^{\pi}(s,a)=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\,\big|\,s_{0}=s,a_{0}=a\right],$ which captures the expected return of taking action $a$ in state $s$, followed by policy $\pi$. Off-policy RL aims to discover the optimal policy $\pi^{*}$ by maximizing the Q-value function using transitions generated from a different behavior policy $\beta$. By decoupling data collection from policy improvement, off-policy methods can achieve high sample efficiency—primarily due to the effective reuse of past transitions stored in a replay buffer. However, when the behavior policy $\beta$ deviates too far from $\pi$, the resulting distributional shift can severely undermine value estimation, a delicate yet critical challenge that continues to motivate much of the recent progress in off-policy learning Duan et al. (2024).
@@ 2.2 Online Planner
Online planning optimizes action sequences at each step by simulating future trajectories under a predictive model, enabling informed action selection Zhan et al. (2024a). This approach typically yields high-quality decisions by leveraging foresight over potential future outcomes. As a result, it is widely adopted in model-based RL for environment interaction, generating higher-quality interaction samples that facilitate more efficient learning and policy improvement.
Among various planning methods, Model Predictive Path Integral (MPPI) is a widely adopted sampling-based planner for continuous control tasks due to its high efficiency Williams et al. (2017). Formally, at each planning step, with the help of a dynamics model, $N_{p}$ action trajectories are sampled from a factorized Gaussian: $a_{t}^{i}\sim\mathcal{N}(\mu_{t},\sigma_{t}^{2}),\quad\text{for }t=0,\dots,H{-}1\text{ and }i=1,\dots,N_{p},$ where $H$ is the predictive horizon. Each trajectory is evaluated using a reward and value model to estimate the future return: $G^{i}=\sum_{t=0}^{H-1}r_{t}^{i}+\hat{v}_{H}^{i},$ where $r_{t}^{i}$ denotes predicted rewards and $\hat{v}_{H}^{i}$ is the terminal value estimate. To update the trajectory distribution, MPPI reweights the $N$ samples using a softmax over the returns: $w^{i}={\exp(G^{i}-\max_{j}G^{j})}/{\sum_{k}\exp(G^{k}-\max_{j}G^{j})}.$ Next, the parameters of the trajectory distribution are updated using weighted statistics: $\mu_{t}\leftarrow\sum_{i}w^{i}a_{t}^{i}.$ This process is repeated over several iterations, gradually converging toward an optimal solution. At test time, the planner executes the first action $\mu_{0}$, while during training, Gaussian noise is added for exploration.
@@ 3.2 Bootstrap Off-policy with World Model
$\mathcal{L}_{\text{align}}=\mathbb{E}_{(s,a,r,s^{\prime})_{0:H}}\sum_{t=0}^{H-1}\sum_{i=1}^{N}w_{i}\left[-\log\pi(a_{i}\mid s_{i})\right].$ (4)
This mechanism ensures that the policy assigns higher probability to high-value actions, effectively guiding it toward promising regions. Besides, since these actions stored in the replay buffer may vary in quality, this value-weighting ensures that the policy prioritizes the most beneficial experiences, thus accelerating learning while accommodating the variability in the planner’s past actions. The Q function here can also be replaced by other critics, such as state value V or advantage A, here we think Q is the most convenient one to access.
@@
$\mathcal{L}_{\text{policy}}=-\mathbb{E}_{(s,a,r,s^{\prime})_{0:H}}\sum_{t=0}^{H-1}\bigg[Q(s,\pi(s))+\lambda_{\text{align}}\cdot\mathcal{L}_{\text{align}}\bigg],$ (5)
@@
The reasons are two-fold. (1) Improved value learning. By aligning the policy with the planner, we reduce the discrepancy between the collected data and the actual policy behavior. This improves the distributional matching during training, enabling the value estimator to learn from more consistent and policy-relevant trajectories. As a result, the predicted values used in planning become more accurate and reliable. (2) Improved representation, reward, and dynamics learning. We adopt a TD-style learning objective that jointly optimizes the value function along with the other components of the world model as shown in (1). As the value predictions become more accurate due to better distributional match, the gradients flowing into the encoder, dynamics model, and reward predictor become more informative, leading to improved overall model quality.
@@
12:  Replay rollout batch $\{(s_{t},a_{\beta},r_{t},s_{t+1})_{t=0}^{H-1}\}\sim\mathcal{D}$
@@ 4.1 Experimental Setup
We choose four representative online RL algorithms as our baselines: (1) SAC Haarnoja et al. (2018): the state-of-the-art model-free off-policy RL algorithm under the maximum entropy framework; (2) DreamerV3 Hafner et al. (2025): the state-of-the-art imagination-driven model-based RL algorithm that learns from imaginary rollouts generated by a world model. To enable a comprehensive comparison, we evaluate DreamerV3 under two different interaction budgets: 2M and 10M iterations; (3) TD-MPC2 Hansen et al. (2023): a planning-driven model-based RL algorithm that performs online planning via an MPPI planner; (4) BMPC Wang et al. (2025): an improved variant of TD-MPC2, where the policy is trained solely to imitate the actions generated by the MPPI with an extra relabeling mechanism to update the historical actions in the replay buffer with the latest planner.
@@
We evaluate our method on a challenging benchmark of 14 high-dimensional locomotion tasks drawn from the DeepMind Control Suite (DMC) Tassa et al. (2018) and the recently proposed Humanoid Bench (H-Bench) Sferrazza et al. (2024). The chosen 7 DMC tasks feature two most complex agents—humanoid (67/21 state/action dims) and dog (223/38)—that demand sophisticated balance and coordination. The other 7 H-Bench tasks raise the difficulty further with long-horizon, goal-directed tasks on the Unitree H1hand robot (151/61), such as walking over slides, traveling over a pole forest without collision, and continuously crossing hurdles. Detailed descriptions are listed in Appendix B.1.
Implementation details. The detailed hyperparameters and reproducibility statement of other baselines are documented in Appendix B.2.
@@ 4.2 Experimental Results
All the training curves are shown in Figure 2 and the detailed numerical results are listed in Table 1. Our method, BOOM, consistently delivers the best Total Average Return (TAR) across all 14 high-dimensional locomotion tasks. These tasks pose significant challenges due to their large state and action spaces, yet BOOM demonstrates remarkable stability and effectiveness.
Results on the DMC Suite.
Our BOOM achieves an average TAR of 877.7, outperforming the previous best BMPC (835.8) by a substantial +5.0%, and exceeding TD-MPC2 by an even larger margin of +17.7%. Notably, BOOM achieves new best results on all humanoid and dog tasks. In Humanoid-run, BOOM outperforms the second-best method by +9.7%, and in Dog-run, it leads by a staggering +21.8%. While SAC and DreamerV3 often fail on these high-dimensional control tasks and achieve near-zero performance, planning-based methods like TD-MPC2 and BMPC perform better but still suffer from instability and limited final returns. In contrast, BOOM learns faster and achieves much stronger final performance across all tasks.
Results on the Humanoid Bench.
Our BOOM again dominates with an average TAR of 820.6. This marks a dramatic +47.7% improvement over DreamerV3 (10M), which is (555.6), and an even more impressive +60.5% gain over BMPC. BOOM sets new records on every single task in the Humanoid Bench. For example, in H1hand-slide, BOOM improves over the second-best method by +110.5%, in H1hand-pole, by +25.8%, and in H1hand-hurdle, by +121.0%, .
@@
Task SAC DreamerV3 ( 2M & 10M iters) TD-MPC2 BMPC BOOM (ours)
@@ 4.3 Ablation Study
We vary the alignment coefficient by testing $0.1\times$ and $10\times$ the default setting ($\lambda_{\text{align}}=\dim(\mathcal{A})/1000$ for DMC and $\dim(\mathcal{A})/50$ for Humanoid-Bench). Results in Figure 3(c) show stable performance across this range, suggesting that BOOM is robust to this hyperparameter and does not require very sensitive tuning.
@@ 5 Related Work
Planning-driven methods use planner rather than policy itself to generate high-quality actions for environment interaction Curi et al. (2020). Early work has shown that solely learning the value and then combining it with the planner can achieve good control performance  Nagabandi et al. (2018); Kocijan et al. (2004); Deisenroth et al. (2013); Lowrey et al. (2019). To further improve performance in high-dimensional tasks, recent approaches have combined online planning with policy learning, where the policy can provide a good initial solution to speed up planningJanner et al. (2019); Wang and Ba (2019); Nguyen et al. (2021); Buckman et al. (2018); Lin et al. (2025). LOOP Sikchi et al. (2022) takes SAC Haarnoja et al. (2018) as the backbone, and employs a planner under the policy behavior constraint for collecting samples. TD-MPC family Hansen et al. (2022, 2023) jointly learns model and value function through TD-learning, achieving strong performance through both algorithmic innovations and implementation advances. It successfully delivers substantial gains over model-free baselines, particularly on complex benchmarks. However, these methods inevitably encounter actor divergence—a mismatch between the planner and the policy. Our approach, BOOM, addresses this challenge by tightly coupling planning and off-policy learning through a bootstrap loop that aligns the policy with the planner’s non-parametric action distribution via a Q-weighted likelihood-free alignment loss, preserving distributional consistency. A recent method BMPC Wang et al. (2025) simplifies the pipeline by discarding explicit policy optimization and directly imitating the planner to avoid actor divergence. However, BMPC ignores the Q-function during training, resulting in lower policy learning efficiency and sensitivity to the variability of historical planner actions in the buffer. This leads to unstable learning, as reflected by its oscillatory training curves in our experiments. We acknowledge a concurrent and close work, TDM(PC)2 Lin et al. (2025), which similarly found that aligning the policy with the planner is beneficial. The major distinction lies in the design of the alignment objective: their approach follows a TD3+BC style using reverse KL, whereas ours is closer to AWAC using critic-guided weights and forward KL.
@@
Imagination-driven methods leverage a learned model to generate synthetic rollouts for policy and value updates Sutton (1991). Modern approaches such as SimPLe Kaiser et al. (2019), IRIS Micheli et al. (2022), IDM Mu et al. (2021), and the Dreamer family Hafner et al. (2019a, 2020, 2023) train latent dynamics models to support actor-critic learning entirely in imagination. These methods offer fast test-time execution and relatively high sample efficiency, but their performance is often limited by compounding model errors over long imagined rollouts. Among them, DreamerV3 Hafner et al. (2025) stands out as a leading representative, demonstrating strong performance across a range of tasks. Unlike planning-based MBRL methods, DreamerV3 interacts by directly sampling actions from its learned policy, typically requiring more iterations to converge. For comprehensive comparison, we evaluate DreamerV3 under 2M and 10M iterations. While it improves with more interaction budget and outperforms TD-MPC2 and BMPC on certain tasks, our BOOM consistently outperforms all baselines across all tasks.
@@
One might notice the similarity between BOOM and offline RL; however, the fundamental paradigms and practical implications of these two settings differ significantly. (1) Learning paradigm. Offline RL centers on a fixed dataset, relying heavily on behavior cloning (BC) to restrict the policy within the dataset’s support Fujimoto et al. (2019). To cautiously improve policy performance, offline RL relaxes BC for poor actions, but always aims to keep the policy close to known data to avoid extrapolation error Prudencio et al. (2023); Ran et al. (2023). BOOM centers on the policy itself as the optimization target. It seeks to maximize Q-values and align with planner-generated actions, both aimed at improving policy performance. Ultimately, the policy and planner co-adapt and converge to the optimal solution through online interaction. (2) Learning objective. In offline RL, maximizing Q-values and BC often conflict Fujimoto and Gu (2021): Q-values outside the dataset distribution cannot be accurately estimated, so maximizing Q risks pushing the policy toward unsupported actions; BC pulls the policy back toward known actions. This tension forces a conservative balance Wang et al. (2022). In BOOM, the two objectives are largely complementary: the planner generally produces higher-quality actions than the policy Zhan et al. (2025). Aligning policy with the planner improves performance and strengthens policy-planner consistency, which in turn leads to more accurate Q estimates. More accurate Q-values then enable better policy improvement and allow the planner to generate higher-quality actions. This positive feedback loop bootstraps policy and planner consistently toward faster convergence to the optimal solution.